# Modelling Conduction Cooling of Superconducting Accelerator Magnets using a Thermal Thin Shell Approximation

### v1

@@ 3.2 TSA performance
The TSA results are also reported in Figs. 56 in black and in Tab. 2 in gray. First, we note that TSA with scaling is closer to the reference model compared to the TSA without scaling, thus highlighting the necessity of introducing the scaling factor. Compared to the fully resolved reference mesh, the TSA overestimates the temperature, while the simulations with a fully resolved mesh are persistently underestimating the reference model.
A quench occurs if the source current in the half-turns surpasses the critical current determined by the half-turn temperature and the local magnetic field. This leads to a rapid increase in temperature. For this reason, employing the TSA for maximum ramp rate estimation is considered a more conservative approach, as it predicts a quench at lower ramp rates. In contrast, the use of a coarse mesh may fail to predict a quench that would occur in the reference.
We conclude that the TSA reaches accuracies below $2\,\%$ before and up to $4\,\%$ after the ramp. This is only slightly worse than the fully resolved mesh with $13$ k DoFs, with errors below $0.8\,\%$ and up to $5\,\%$, but better than the $6$ k DoFs model. Regarding the computational cost in Tab. 2, we can see a reduction in total generation and solving time for the TSA model. It should be noted that the total computation time for every model includes the EM system, which requires about 72 s for generation and 15 s for solving. As generation wall time scales approximately linearly with the number of DoFs, while the solving wall time scales superlinearly, the TSA leads to significantly faster simulations. The TSA yields a total speed-up between $250\%$ and $517\%$, compared to the models marked in bold in Tab. 2 and meshes shown in Fig. 7. Although this speed-up is less significant than for quench simulations [12], where fine meshes in the insulation layer were needed due to the large temperature gradients and large heat fluxes, it still shows the computational efficiency of the TSA for the ramping simulations conducted in this work.
Note that in 2D simulations, generation times can become larger than solution times. In this context, a dedicated assembly procedure has been implemented in GetDP 4.0 [8] to efficiently handle the large number of terms in the weak formulations associated with each TSL. In particular, the current model description may lead to some TSA meshes having more weak formulation terms than DoFs. Special care is taken to perform basic elementary operations, such as element-support mapping, just once during the very first assembly. This approach avoids the significant overhead otherwise observed during subsequent assemblies for these application-specific edge cases, as it reduced the generation time reported in Tab. 2 by a factor of 9 for the 3 k DoFs model.
@@ 3.3 Example: cooling hole position study
The developed model is next employed to assess cooling strategies in the collar. Only the TSA models are used since fully meshed models are too costly and the TSA is sufficiently accurate, as was shown in Sec. 3.2. Cooling is implemented via the Robin boundary condition (3) at the boundaries of the cooling holes, denoted as $\Gamma_{\mathrm{Cool}}$ in Fig. 1.
Several configurations, varying in the number and positions of the cooling holes, are considered. Configuration $A$ compromises eight cooling holes, as shown in Fig. 8. Configuration $B$ also consist of eight cooling holes, but positioned closer to the half-turns. Fig. 9 shows the hole numbering in configuration $A$, with $A_{o}$ and $A_{e}$ indicating odd- and even-numbered holes when only half are active. The two-hole configuration $A_{1,7}$, equivalent to the $A_{3,5}$ by symmetry, is selected to test two-hole cooling.
Table 3: Lower and upper bounds on the maximal ramp rate for a non-quenching simulation, considering different cooling configurations (Figs. 8 and 9) and collar materials. The increase in collar temperature with respect to the initial temperature of $1.9$ K is reported for the lower (i.e. non-quenching) simulations, and its temperature distribution is visualised in Fig. 10 for the $A_{o}$ configuration.
In order to find the fastest ramp rate without thermal runaway, an upper and lower bounds is estimated by varying the rates until a quenching and non-quenching simulation is found. For the non-quenching simulation, the increase in collar temperature is reported. The ramp rate interval size is fixed with a difference of $1$ s ramp time.
Concerning the aluminium alloy collar, the ramp rates reported in Tab. 3 are roughly identical for all cooling positions; with configurations $A_{e}$ and $A_{1,7}$ requiring the slowest ramp rate. Looking at the increase in collar temperature, which is much smaller than for stainless steel, configuration $A$ is able to keep the collar the coldest. Note that configuration $A_{o}$ only uses 4 cooling holes, while still performing only slightly worse than configuration $A$. To conclude, for the aluminium alloy collar the cooling positions are of little importance, and the use of only two cooling holes ($A_{1,7}$) may be considered.
Acceptable ramp rates for the stainless steel collar are much lower, which is expected due to the reduced thermal conductivity compared to aluminium alloy. As the heat distributes much slower, the positioning of the cooling holes is more crucial. The heat capacity of steel is also lower, which means that it requires less heat to increase its temperature. This is visible from the increase in collar temperature, which can be explained by these effects. Configuration $A$ shows a small advantage over configuration $B$, allowing for a slightly faster ramp. Configurations $A_{o}$ and $A_{e}$ are comparable, but considerably slower than cooling with 8 cooling holes. This suggests that reducing the number of cooling holes to 4 or less would probably no longer be beneficial, since the cooling rates would be reduced significantly.
@@
Figure 9: The numbering for the cooling holes in configuration $A$. In configuration $A_{o}$ ($A_{e}$), only the odd (even) numbered cooling holes in orange (blue) are active, while in $A_{1,7}$, cooling is provided through holes 1 and 7.
Figure 10: Temperature distribution (in K) for the $A_{o}$ cooling configuration with an aluminium alloy collar and ramp rate $110.52$ A/s (left), and a stainless steel collar and ramp rate $67.09$ A/s (right), each at the time (in s) when $\Delta T_{\mathrm{col}}$ in Tab. 3 is reached.
@@ 4 Conclusion
To achieve this, we generalised the TSA to the collar and pole interfaces, thereby avoiding the need for explicit meshing of the insulation layers. This increases meshing flexibility, while maintaining an acceptable level of accuracy. In this work, a necessary length correction was introduced to account for the tangential heat flux between neighbouring thin shell layers, and applied to a 12 T dipole magnet as a use case. Tangential and radial contributions treated separately, such that the correction can be generalised to other magnets. Additionally, an automated workflow for the TSA is implemented, which would be substantially more difficult to construct in the case of the fully resolved mesh models.
The TSA models were shown to reproduce the temperature evolution of the fully resolved mesh reference model of a 12 T dipole magnet within a relative error of approximately 2–4 %, with a computational speed-up up to 5 times. Although the TSA does not converge towards the reference for finer mesh discretisations, it provides a robust and efficient alternative for estimating maximum allowable ramp rates with sufficient accuracy. Furthermore, the method remains applicable under quenching or steady-state conditions.
As an illustration, the TSA model was used to evaluate cooling strategies for the collar region. Aluminium alloy collars were largely insensitive to the cooling hole configuration, unlike stainless steel collars. This indicates that an aluminium alloy collar can improve thermal performance, allowing for similar current ramp rates with fewer cooling holes.

### v2

@@ 3.2 TSA performance
From Fig. 5 we note that TSA with scaling is closer to the reference model compared to the TSA without scaling, thus highlighting the necessity of introducing the scaling factor. Compared to the fully resolved reference mesh in yellow, Fig. 5 also shows that the TSA overestimates the temperature, while the simulations with a fully resolved mesh are persistently underestimating the reference model.
A quench occurs if the source current in the half-turns surpasses the critical current determined by the half-turn temperature and the local magnetic field. This leads to a rapid increase in temperature. For this reason, employing the TSA for maximum ramp rate estimation is considered a more conservative approach, as it predicts a quench at lower ramp rates. In contrast, the use of a coarse mesh may fail to predict a quench.
We conclude that the TSA reaches accuracies below $2\,\%$ during and below $4\,\%$ after the ramp. This is slightly worse than the fully resolved mesh with $13$ k DoFs, with errors below $0.8\,\%$ and up to $5\,\%$, but better than the $6$ k DoFs model. Regarding the computational cost in Tab. 2, we can see a reduction in total generation and solving time for the TSA model. It should be noted that the total computation time for every model includes the EM system, which requires about 72 s for generation and 15 s for solving. As wall generation time scales approximately linearly with the number of DoFs, while the wall solving time scales superlinearly, the TSA leads to significantly faster simulations. The TSA yields a total speed-up between $250\%$ and $517\%$, compared to the models marked in bold in Tab. 2 and meshes shown in Fig. 7. Although the observed speed-up in this work is smaller than for quench simulations reported in [12], this difference is expected. In [12], simulations require significantly finer meshes in the insulation layer to resolve steep temperature gradients and large heat fluxes. Nevertheless, the TSA still achieves a noticeable speed-up for the ramping simulations considered here.
In 2D simulations, generation times can become larger than solution times. In this context, a dedicated assembly procedure has been implemented in GetDP 4.0 [8] to efficiently handle the large number of terms in the weak formulations associated with each TSL. In particular, the current model description may lead to some TSA meshes having more weak formulation terms than DoFs. Special care is taken to perform basic elementary operations, such as element-support mapping, just once during the very first assembly. This approach avoids the significant overhead otherwise observed during subsequent assemblies for these application-specific edge cases, as it reduced the generation time reported in Tab. 2 by a factor of 9 for the 3 k DoFs model.
@@ 3.3 Example: cooling hole position study
The developed model is now employed to assess cooling strategies in the collar. Only the TSA models are used since fully meshed models are too costly. Cooling is implemented via the Robin boundary condition (3) at the boundaries of the cooling holes, denoted as $\Gamma_{\mathrm{Cool}}$ in Fig. 1.
Several configurations, varying in the number and positions of the cooling holes, are considered. Configuration $A$ compromises eight cooling holes, as shown in Fig. 8, while configuration $B$ also consist of eight cooling holes, but positioned closer to the half-turns. Fig. 9 shows the hole numbering in configuration $A$, with $A_{o}$ and $A_{e}$ indicating that only odd- and even-numbered holes are active respectively. The two-hole configuration $A_{1,7}$, equivalent to $A_{3,5}$ by symmetry, is selected to test two-hole cooling.
Table 3: Lower and upper bounds on the maximal ramp rate considering different cooling configurations (Figs. 8 and 9) and collar materials. The increase in collar temperature with respect to the initial temperature of $1.9$ K is reported for the lower bound (i.e. non-quenching) simulations.
In order to find the fastest ramp rate without thermal runaway, an upper and lower bounds are estimated by varying the rates until a quenching and non-quenching simulation is found. For the non-quenching simulation, the increase in collar temperature is reported. The ramp time is discretized in fixed increments of $1$ s, the corresponding ramp rates are reported as ranges.
Concerning the aluminium alloy collar, the ramp rates reported in Tab. 3 are roughly identical for all cooling positions; with configurations $A_{e}$ and $A_{1,7}$ requiring the slowest ramp rate. Looking at the increase in collar temperature, which is much smaller than for stainless steel, configuration $A$ is able to keep the collar the coldest. Note that configuration $A_{o}$ only uses 4 cooling holes, but performs only slightly worse than configuration $A$. To conclude, for the aluminium alloy collar the cooling positions are of little importance, and the use of only two cooling holes ($A_{1,7}$) may be considered.
Acceptable ramp rates for the stainless steel collar are much lower, which is expected due to the reduced thermal conductivity compared to aluminium alloy. As the heat distributes much slower, the positioning of the cooling holes is more crucial. The heat capacity of steel is also lower, which means that it requires less heat to increase its temperature. This is visible from the increase in collar temperature, which can be explained by these effects. Configuration $A$ shows a small advantage over configuration $B$, allowing for a slightly faster ramp. Configurations $A_{o}$ and $A_{e}$ are comparable, but considerably slower than cooling with 8 cooling holes.
@@
Figure 9: The numbering for the cooling holes in configuration $A$. In configuration $A_{o}$ ($A_{e}$), only the odd (even) numbered cooling holes are active, while in $A_{1,7}$, cooling is provided through holes 1 and 7.
Figure 10: Temperature distribution (in K) for the $A_{o}$ cooling configuration with an aluminium alloy collar and ramp rate of $110.52$ A/s (left), and a stainless steel collar and ramp rate of $67.09$ A/s (right), each at the time (in s) when $\Delta T_{\mathrm{col}}$ in Tab. 3 is reached.
@@ 4 Conclusion
To achieve this, we generalised the TSA to the collar and pole interfaces, thereby avoiding the need for explicit meshing of the insulation layers. This increases meshing flexibility, while maintaining an acceptable level of accuracy. In this work, a necessary length correction was introduced to account for the tangential heat flux between neighbouring thin shell layers. This was then applied to the simulation of a 12 T dipole magnet as a use case. Tangential and radial contributions are treated separately, such that the correction can be generalised to other magnets. Additionally, an automated workflow for the TSA is implemented, which would be substantially more difficult to construct in the case of fully resolved mesh models.
The TSA models were shown to reproduce the temperature evolution of the fully resolved mesh reference model of a 12 T dipole magnet within a relative error of approximately 2–4 %. The computational time required is nevertheless reduced by up to a factor 5. Although the TSA does not converge towards the reference for finer mesh discretisations, it provides a robust and efficient alternative for estimating maximum allowable ramp rates. Furthermore, the method remains applicable under both quenching and steady-state conditions.
As an illustration, the TSA model was used to evaluate cooling strategies for the collar region. It was shown that aluminium alloy collars were largely insensitive to the cooling hole configuration, unlike stainless steel collars. This shows that aluminium alloy collars can improve thermal performance, allowing for similar current ramp rates with fewer cooling holes.